A particle moves along a path y=2cos 3x in such a way that the component of it velocity in x direction is 4 units. find its acceleration
1
Expert's answer
2019-12-02T09:24:33-0500

Given path,

"y=2cos 3x"


Differentiate with respect to x, We get


"\frac {dy}{dx} = 2 \times ( - sin 3x ) \times 3"



"\frac {dy}{dx} = -6 \times sin 3x"

But, We have

"\frac {dy}{dx} = 4"


So,

"-6 \times sin 3x = 4"


"sin 3x = - \frac {2}{3}"

We know ,

"sin^2 3x + cos^2 3x = 1"


"cos^2 3x = 1 - sin^2 3x"


"cos^2 3x = 1 - (\frac {-2}{3})^2 = 1 - \frac {4}{9} = \frac {5}{9}"

"cos 3x = \frac {\sqrt 5}{3}"


"Acceleration =\frac {d^2 y} {dx^2}"


"= \frac {d}{dx} ( - 6 \times sin 3x)"


"- 6 \times cos 3x \times 3 = - 18 \times cos 3x = - 18 \times \frac {\sqrt 5}{3} = - 6 \times {\sqrt 5}"

Answer:

Acceleration = "- 6 \times \sqrt 5"